# Hyperparameter Optimization in Machine Learning

### v1

@@ 6.2 Particle swarm optimization
Assume for simplicity that $\Lambda=\mathbb{R}^{m}$, and let $\{(\lambda_{1}^{j},v_{1}^{j})\}_{j=1}^{N}$ be an initial population of hyperparameters and their velocities, both drawn randomly from suitable distributions (e.g. uniform distributions). lorenzo2017particle propose an application of PSO to hyperparameter optimization with the following equation:
$\lambda_{t+1}^{j}$ $=\lambda_{t}^{j}+v^{t}_{j}$
$v_{t+1}^{j}$ $=\alpha_{1}v_{t}+\alpha_{2}r_{l}(\bar{\lambda}^{j}-\lambda_{t})+\alpha_{2}r_{g}(\lambda^{*}-\lambda_{t})$
where $\alpha_{1},\alpha_{2},\alpha_{3}\in\mathbb{R}^{+}$ are three coefficients, $r_{l},r_{g}\in\mathbb{R}^{m}$ are two randomly drawn vectors and $\bar{\lambda}^{j}$ and $\lambda^{*}$ are the particle’s best and population best points found so far, respectively. The second term in the velocity update equation can be interpreted as a local memory attractor, while the third term is a global attractor. The randomness introduced by $r_{l}$ and $r_{g}$ serves to encourage diversity throughout the search.
@@ 6.3 Discussion
In some cases, the response function, $f(\lambda)$, is differentiable with respect to $\lambda$, enabling the application of optimization algorithms that require the gradients of their objective. Of course assuming that $f(\lambda)$ be differentiable excludes discrete hyperparameters such as discrete choices (for example, which nonlinear activation function should be used in a layer of a neural network) or integer-valued hyperparameters (such as the dimension of a layer). There is, however, a distinct advantage of gradient-based methods for HPO, namely the possibility of tuning a (very) high-dimensional hyperparameter vector $\lambda$, which would be rather difficult to do with model-based algorithms. This opportunity allows us to completely rethink when a variable should be considered as a parameter or as a hyperparameter. For example, [maclaurin2015gradient] suggested to consider the initial weights of a neural network as a large hyperparameter vector, an approach that in their own words effectively “blurs the distinction between learning and meta-learning.” Indeed, gradient-based methods reveal some deep connections between HPO and some forms of meta-learning (see also Section 8.4). For example for example, the idea of tuning initial weights from several learning tasks was shortly after successfully exploited in [finn_model-agnostic_2017] in their model-agnostic meta learner.
@@ 7.2 Hypergradients via implicit differentiation
$A(\mathcal{D},\lambda)\in\operatorname*{arg\,min}_{u\in\mathcal{W}}L(u,\lambda),$ (7.3)
where $L:\mathcal{W}\times\Lambda\to\mathbb{R}$, with $\mathcal{W}\subseteq\mathbb{R}^{d}$ is an objective function (the training error) that depends on the data.  [^1] We leave such dependence implicit for improving readability. If $L$ is a smooth function w.r.t. $u$ then, the first-order optimality condition implies that
$\nabla_{u}L(u,\lambda)=0.$ (7.4)
Notably, neural networks and many related deep learning techniques are formulated in these terms. If $L$ is twice differentiable w.r.t. $u$ at a point $\bar{\lambda}\in\Lambda$ and $\partial_{w}^{2}L\in\mathbb{R}^{d\times d}$ is invertible, we can invoke the implicit function theorem [krantz2012implicit] to (7.4) which assures us that (locally) there exist a function $w:U\subset\Lambda\to\mathcal{W}$ such that $\nabla_{u}L(u,\lambda)_{|u=w(\lambda)}=0$ for all $\lambda\in U$, and $A(\mathcal{D},\lambda)=w(\lambda)$. In such case, we can locally express the HPO objective as the nested function $f(\lambda)=\mathscr{M}(w(\lambda),\mathcal{V})$. Now, if $\mathscr{M}$ is differentiable, we have that
@@
$\partial_{\lambda}w(\lambda)=-\left[\partial_{u}^{2}L(u,\lambda)_{|u=w(\lambda)}\right]^{-1}\partial^{2}_{u,\lambda}L(u,\lambda)_{|u=w(\lambda)}.$ (7.6)
@@
Let us denote by $w_{T}$ an approximate solution of the inner problem, e.g. found with an iterative optimization algorithm. Then, after substituting $w(\lambda)$ with $w_{T}$ and plugging (7.6) into (7.5), one can set $q^{\intercal}=-\partial_{u}\mathscr{M}\left[\partial_{u}^{2}L\right]^{-1}\in\mathds{R}^{d}$ and approximately solve the linear system
$\partial^{2}_{u}L(u,\lambda)_{|u=w_{T}}q=-\nabla_{u}\mathscr{M}(u)_{|u=w_{T}}.$ (7.7)
@@
$\nabla_{T,K}f(\lambda)=\left[\partial^{2}_{u,\lambda}L(u,\lambda)_{|u=w_{T}}\right]^{\intercal}q_{T,K}.$ (7.8)
@@ 7.3 Hypergradients via iterative differentiation
Methods based on iterative differentiation further assume that solutions to the learning algorithm problem (7.3) are sought (with approximation) by iterating an optimization dynamics for $T$ steps:
@@
We call $s_{t}=[w^{\mathsf{T}}_{t},v^{\mathsf{T}}_{t}]^{\mathsf{T}}\in\mathds{R}^{d}$ the “state” of the optimizer, which contains the current parameters, $w_{t}$ and a (possibly empty) set of auxiliary variables the $v_{t}$, such as velocities and moments for algorithms such as Adam and RMSProp. This process yields a solution $w_{T}=w_{T}(\lambda)$. The transition map $\Phi:(\mathds{R}^{d}\times\mathds{R}^{m}\times\mathscr{D})\rightarrow\mathds{R}^{d}$ is assumed to be differentiable and describes the actual operations associated with a specific optimizer. Typically, $\Phi_{t}$ depends explicitly on the current minibatch $\mathcal{D}_{t}$ (e.g. for stochastic gradient descent) and it internally requires the computation of the inner objective gradients
$g_{t}\doteq\nabla_{w}\bigl{.}L(w,\lambda,\mathcal{D}_{t})\bigr{|}_{w=w_{t-1}}.$
@@ 7.4 Fixed-point methods
which could be, for instance, gradient descent $\Phi(w,\lambda)=w-\eta\nabla L$.  [^3] As in the iterative case, the dynamics may be defined on an augmented state, for example when using gradient descent with momentum. Then, the application of the implicit function theorem to (7.15) yields to
@@ 7.5 Discussion
Implicit differentiation methods are “oblivious” to the way the approximate solution $w_{T}$ has been found. This means that they have a computational advantage in memory over the reverse-mode differentiation as they do not require storing the entire dynamics. If implemented using algorithmic differentiation, they have a run-time advantage over forward-mode differentiation since they only deal with vector quantities (the $q_{T,K}$’s). However, for this very reason, they do not directly allow computing the gradient of hyperparameters related to the optimization of the inner problem, such as the initialization mapping or the learning rate. Furthermore, they are also quite sensitive to the proprieties of the inner problem (e.g. for the fixed-point method, the dynamics must be a contraction), and therefore are applicable to a narrower class of problems.
[maybe move and shorten] Computational disadvantages of plain iterative reverse mode are related to memory requirements. Indeed, the memory complexity is $O(TW)$, with $W$ the number of weights, which can quickly become prohibitive when tuning the hyperparameters associated with modern networks. The problem has been addressed in different ways. maclaurin2015gradient suggested to avoid completely the storage of $s_{t}$, recomputing it in the backward pass by reversing the learning dynamics (7.9). Since the operation is numerically unstable, reconstructed initial weights may be very different from the original ones. As a remedy, they proposed an effective numerical strategy capable of overcoming the loss of information due to finite precision arithmetic. A general technique for reducing the memory requirements in reverse mode AD is checkpointing [Griewank08:_evaluat_deriv, Ch. 12], where only a subset of the possible checkpoints is stored, namely $w_{jC}$ for $j=1,\dots\lceil T/C\rceil$, being $C$ a given window size. In so doing, one must redo the forward pass from $jC$ to $(j+1)C-1$ to recover the missing $w_{t}$, effectively doubling the computational effort. It is easy to see that the best window size is $O(\sqrt{T})$, which allows us to complete hypergradient calculations with a memory size of $O(W\sqrt{T})$. This technique has been considered in [shaban2019truncated], together with a more radical truncation strategy (similar to truncated backpropagation through time [DBLP:journals/neco/WilliamsP90]), which however does not recover exact hypergradients.
@@
In the last six chapters we covered the main classes of algorithms for hyperparamter optimization. Here, moving away from particular techniques, we discuss horizontally advanced topics in HPO such as multi-objective and constrained optimization, online HPO and hyperparameter schedules and neural architecture search. We conclude the chapter with a section on meta-learning, a closely related topic in the literature, discussing similarities and differences with the typical hyperparameter optimization setting.
@@ 8.1 Optimizing for multiple objectives
In many real world settings there is not just one, but multiple, potentially conflicting, objectives of interest. For example when working with hardware-constraint systems one desires predictive models that are accurate and have low resource consumption. Assume one needs to decide between two solutions $\lambda_{1},\lambda_{2}$. Each individual objective induces a distinct ordinal relationship between the candidates, but while $\lambda_{1}$ might dominate $\lambda_{2}$ in one objective it might be worse in another. Because there is rarely a single best solution, the multi-objective hyperparameter optimization (MO) problem is to identify the Pareto front – the set of all non-dominated solutions – to help users take an informed trade-off decision.
More formally, consider a function $f:\Lambda\to\mathcal{Y}$ mapping points in domain $\Lambda$ to a space $\mathcal{Y}$, which in the following we consider to be a $n$-dimensional objective space that we aim to minimize. For two points $\lambda_{1},\lambda_{2}\in\Lambda$, we say $\lambda\_1$ weakly dominates $\lambda_{2}$ if $f(\lambda_{1})_{i}\leq f(\lambda_{2})_{i}$ for all objectives $i\in[n]$. We write $\lambda_{1}\succeq\lambda_{2}$ to denote weak domination.
Furthermore, $\lambda_{1}$ is said to dominate $\lambda_{2}$ if $\lambda_{1}\succeq\lambda_{2}$ and $f(\lambda_{1})_{i}<f(\lambda_{2})_{i}$ for some objective $i$. We write $\lambda_{1}\succ\lambda_{2}$ for domination. The Pareto front $\mathcal{P}_{f}$ contains all nondominated points in $\Lambda$, formally: $\mathcal{P}_{f}=\{\lambda\in\Lambda\,|\,\nexists\,\lambda^{\prime}\in\Lambda:\lambda^{\prime}\succ\lambda\}$.
Since the Pareto front is often infinite, multi-objective optimizers seek a finite approximation set $A\subset\Lambda$ of objective vectors near the Pareto front. The quality of $A$ can be measured by the dominated hypervolume indicator $\mathcal{H}$ [Zitzler1998:Multiobjective]. Given a reference point $\mathbf{r}$, $\mathcal{H}(A)$ equals the hypervolume dominated by $A$:
$\mathcal{H}(A)=\text{Vol}\left(\{\lambda\in\mathcal{Y}=\mathbb{R}^{n}\,|\,\exists\,\lambda^{\prime}\in A:\lambda^{\prime}\succeq\lambda\wedge\lambda\succeq\mathbf{r}\}\right).$
Computing $\mathcal{H}(A)$ involves partitioning the space into hypercubes, an operation that scales exponentially with the number of objectives $n$. This exponential scaling poses a key bottleneck for hypervolume-based multi-objective optimizers.
@@
A popular example of MO scenario is the case of algorithmic fairness which aims to train accurate models subject to fairness criteria. Prior work in the field has proposed several fairness criteria, whose suitability depends on the application domain, and it has been shown that criteria, commonly referred to acceptance-based and error-based, can be in conflict with each other [friedler2016possibility, fairness2018verma, kleinberg2018inherent]. However, obtaining solutions that satisfy various definitions to some (albeit) imperfect extent and expose empirical trade-offs between various definitions can still be important from societal perspectives.
Another way to tackle MO scenarios is the case of gradient-free optimization with stochastic constraints. For example, when enhancing the accuracy of a machine learning model, considerations may include constraints on model memory or prediction latency. This occurs in the practical application of deploying machine learning models on mobile devices, where memory cannot exceed a certain pre-specified threshold, or when models power devices with tight requirements on inference latency. The objective and constraint functions in such cases are real-valued and observed together. Alternatively, in tuning a deep neural network (DNN), one may aim to prevent training failures due to out-of-memory errors. Here, the constraint feedback is binary, indicating whether the constraint is violated, and the objective is not observed under constraint violation. It is common to treat constraints with equal importance to objectives, employing a joint or conditionally independent random function model. Unfeasible evaluations incur a cost, such as wasted resources or compute node failure, necessitating minimization of their occurrence. However, avoiding the unfeasible region should not compromise convergence to optimal hyperparameters. Often, optimal configurations lie at the boundary of the feasible region, presenting a challenge in balancing the costs of unfeasible evaluations and the need for exploration.

### v2

@@ 6.4 Particle swarm optimization
Assume for simplicity that $\Lambda=\mathbb{R}^{m}$, and let $\{(\lambda_{1}^{j},v_{1}^{j})\}_{j=1}^{N}$ be an initial population of hyperparameter and velocity pairs, both drawn randomly from suitable distributions (e.g. uniform distributions). lorenzo2017particle propose an application of PSO to hyperparameter optimization that uses the following equation:
$\lambda_{k+1}^{j}$ $=\lambda_{k}^{j}+v_{k}^{j}$ (6.2)
$v_{k+1}^{j}$ $=\alpha_{1}v_{k}+\alpha_{2}r_{l}\circ(\bar{\lambda}_{k}^{j}-\lambda_{k})+\alpha_{2}r_{g}\circ(\lambda^{*}-\lambda_{k})$ (6.3)
where $\alpha_{1},\alpha_{2},\alpha_{3}\in\mathbb{R}^{+}$ are three coefficients, $r_{l},r_{g}\in\mathbb{R}^{m}$ are two randomly drawn vectors and $\circ$ is the Hadamard product. The configuration $\bar{\lambda}_{k}^{j}$ and $\lambda_{k}^{*}$ are, respectively, the particle’s best and population best points found so far, that is,
$\bar{\lambda}_{k}^{j}=\operatorname*{arg\,min}_{\lambda\in\bar{\Lambda}^{j}_{k}}f(\lambda)\quad\text{for }j\in[N],$
where $\bar{\Lambda}^{j}_{k}=\left\{\lambda^{j}_{t}\,:\,t\in[k-1]\right\}$, and
$\lambda^{*}_{k}=\operatorname*{arg\,min}_{\lambda\in\Lambda^{*}_{k}}f(\lambda),$
where $\Lambda^{*}_{k}=\left\{\lambda_{t}^{j}\,:\,t\in[k-1],\,j\in[N]\right\}$. The second term in the velocity update of Eq. (6.3) can be interpreted as a local memory attractor that pushes the current configuration toward the individual best configuration found so far, while the third term is a global attractor that pushes toward the global best configuration found so far. The randomness introduced by $r_{l}$ and $r_{g}$ serves to encourage diversity throughout the search. Referring to Algorithm 3, each individual distribution is implicitly modelled by the Equations (6.2)-(6.3).
@@ 6.5 Discussion
In some cases, the response function, $f(\lambda)$, is differentiable with respect to $\lambda$, enabling the application of optimization algorithms that require the gradients of their objective. Assuming that $f(\lambda)$ be differentiable excludes the direct optimization of discrete hyperparameters such as discrete choices (for example, which nonlinear activation function should be used in a layer of a neural network) or integer-valued hyperparameters (such as the dimension of a layer). There is, however, a distinct advantage of gradient-based methods for HPO, namely the possibility of tuning a (very) high-dimensional hyperparameter vector $\lambda$, which would be rather difficult to do with model-based algorithms. This opportunity allows us to completely rethink when a variable should be considered as a parameter or as a hyperparameter. For example, maclaurin2015gradient suggested to consider the initial weights of a neural network as a large hyperparameter vector, an approach that in their own words effectively “blurs the distinction between learning and meta-learning.” Indeed, gradient-based methods reveal some deep connections between HPO and some forms of meta-learning, which we expand on in Section 8.4). For example, the idea of tuning initial weights from several learning tasks was shortly after successfully exploited in [finn_model-agnostic_2017] in their model-agnostic meta learner.
The crux of gradient-based HPO lies in computing approximate gradients of the response function w.r.t. hyperparameters. In this chapter, we cover three main technical approaches: implicit differentiation (Section 7.2), iterative differentiation (Section 7.3) and fixed-point iterations (Section 7.4). Next, we explore some use cases in Section 7.5. We touch on extensions that allow for the gradient-based optimization of discrete hyperparameters in Section 7.6, and conclude with a discussion centred on computational considerations in Section 7.7.
@@ 7.2 Hypergradients via implicit differentiation
$A(\mathcal{D},\lambda)\in\operatorname*{arg\,min}_{u\in\mathcal{W}}\mathcal{J}(u,\lambda),$ (7.3)
where $\mathcal{J}:\mathcal{W}\times\Lambda\to\mathbb{R}$, with $\mathcal{W}\subseteq\mathbb{R}^{d}$ is an objective function (the training error) that depends on the data – we have seen a concrete example in (2.3).  [^1] We leave dependence on data implicit for improving readability. If $\mathcal{J}$ is a smooth function w.r.t. $u$ then, the first-order optimality condition implies that
$\nabla_{u}\mathcal{J}(u,\lambda)=0.$ (7.4)
Notably, neural networks and many related deep learning techniques are formulated in these terms. If $\mathcal{J}$ is twice differentiable w.r.t. $u$ at a point $\bar{\lambda}\in\Lambda$ and $\partial_{w}^{2}\mathcal{J}\in\mathbb{R}^{d\times d}$ is invertible, we can invoke the implicit function theorem [krantz2012implicit] to (7.4) which assures us that (locally) there exist a function $w:U\subset\Lambda\to\mathcal{W}$ such that $\nabla_{u}\mathcal{J}(u,\lambda)_{|u=w(\lambda)}=0$ for all $\lambda\in U$, and $A(\mathcal{D},\lambda)=w(\lambda)$. In such case, we can locally express the HPO objective as the nested function $f(\lambda)=\mathscr{M}(w(\lambda),\mathcal{V})$. Now, if $\mathscr{M}$ is differentiable, we have that
@@
$\partial_{\lambda}w(\lambda)=-\left[\partial_{u}^{2}\mathcal{J}(u,\lambda)_{|u=w(\lambda)}\right]^{-1}\partial^{2}_{u,\lambda}\mathcal{J}(u,\lambda)_{|u=w(\lambda)}.$ (7.6)
@@
Let us denote by $w_{T}$ an approximate solution of the inner problem, e.g. found with an iterative optimization algorithm. Then, after substituting $w(\lambda)$ with $w_{T}$ and plugging (7.6) into (7.5), one can set $q^{\intercal}=-\partial_{u}\mathscr{M}\left[\partial_{u}^{2}\mathcal{J}\right]^{-1}\in\mathds{R}^{d}$ and approximately solve the linear system
$\partial^{2}_{u}\mathcal{J}(u,\lambda)_{|u=w_{T}}q=-\nabla_{u}\mathscr{M}(u)_{|u=w_{T}}.$ (7.7)
@@
$\nabla_{T,K}f(\lambda)=\left[\partial^{2}_{u,\lambda}\mathcal{J}(u,\lambda)_{|u=w_{T}}\right]^{\intercal}q_{T,K}.$ (7.8)
@@ 7.3 Hypergradients via iterative differentiation
Another approach to computing hypergradients involves explicitly accounting for an iterative training dynamics. Methods based on iterative differentiation further assume that solutions to the learning algorithm problem (7.3) are sought (with approximation) by iterating an optimization dynamics for $T$ steps:
@@
We call $s_{t}=[w^{\mathsf{T}}_{t},v^{\mathsf{T}}_{t}]^{\mathsf{T}}\in\mathds{R}^{d}$ the “state” of the optimizer, which contains the current parameters, $w_{t}$ and a (possibly empty) set of auxiliary variables the $v_{t}$ – e.g. velocities and moments for algorithms such as Adam and RMSProp. This process yields a solution $w_{T}=w_{T}(\lambda)$. The transition maps $\Phi_{t}:(\mathds{R}^{d}\times\mathds{R}^{m}\times\mathscr{D})\rightarrow\mathds{R}^{d}$ are assumed to be differentiable and describe the actual operations associated with a specific optimizer. Typically, $\Phi_{t}$ depends explicitly on the current minibatch $\mathcal{D}_{t}$ (e.g. for stochastic gradient descent) and it internally requires the computation of the inner objective gradients
$g_{t}\doteq\nabla_{w}\bigl{.}\mathcal{J}(w,\lambda,\mathcal{D}_{t})\bigr{|}_{w=w_{t-1}}.$
@@ 7.4 Fixed-point methods
which could be, for instance, gradient descent $\Phi(w,\lambda)=w-\eta\nabla\mathcal{J}$.  [^3] As in the iterative case, the dynamics may be defined on an augmented state, for example when using gradient descent with momentum. Then, the application of the implicit function theorem to (7.15) yields to
@@
7.5 Use cases
Gradient-based approaches have mostly found application in academic contexts. Their flexibility, however, allows us to tune very high-dimensional hyperparameter vectors, as showcased in several papers.
One possibility is to attach a hyperparameter to every weight in a deep network. We have already mentioned the idea proposed by maclaurin2015gradient of regarding the initial values for the weights as hyperparameters. In the same paper, the authors also suggest a closely related opportunity where a regularization penalty is associated to every individual weights. The meta-learning technique presented in [franceschi2018bilevel] can be interpreted as treating all weights in the first layers of a deep network as hyperparameters, and optimize the loss on new tasks with respect to the weights in the classification head.
Another direction that has been studied is to attach hyperparameters to data. maclaurin2015gradient demonstrated an extreme experiments where image pixels themselves are hyperparameters. In a more realistic scenario, hypergradients have been used to tune the hyperparameters of differentiable data augmentation procedures, yielding significantly speedups without performance loss compared to other data augmentation searching methods [lin:2019:online_hyperparameter_learning, hataya:2020:faster_autoaugment_learning]. A different opportunity is to assign a cost weight to each example in a supervised learning setting so that the overall loss for a training set of $n$ examples is $\sum_{i=1}^{n}\lambda^{(i)}\mathcal{J}(h(x^{(i)}),y^{(i)})$. Here we have $n$ hyperparameters $\lambda^{(i)}$. franceschi2017forward show how this idea can be exploited to filter out from the training set examples with noisy (or perhaps poisoned) labels. In this regards, it is interesting to note that (although not directly related to HPO) the problem of building training set attacks on linear models can be formulated in terms of bilevel optimization [mei:2015:using_machine_teaching] and has commonalities with gradient-based HPO in its mathematical structure.
We finally note here that tuning the whole learning rate schedule has also been tackled with hypergradient approaches [baydin_2017_online, donini2019opt, micaelli2021gradient, sharifnassab:2024:metaoptimize_framework_optimizing]. This specific problem is covered in more detail in Section 8.2.
7.6 Probabilistic reparameterization of discrete hyperparameters
The presence of discrete hyperparameters into the response function – and therefore into the training objective or dynamics – prevents a straightforward application of gradient-based HPO techniques. For some such hyperparameters, however, it might make sense to reformulate them as random variable that follow some discrete distribution. For instance, consider a discrete binary choice $\lambda\in\{0,1\}$, e.g. indicating whether to use an $L^{1}$ or $L^{2}$ regularizer in (2.3). Then, one may define a discrete binary random variable $\iota$ as follows:
$\iota\sim\mathrm{Ber}(\theta),\qquad\theta\in[0,1]$ (7.18)
and reinterpret $\lambda$ as a drawing of $\iota$, where $\mathrm{Ber}(\theta)$ is the Bernoulli distribution with success parameter $\theta$. 444 Similarly, one may reformulate hyperparmeters encoding multiple choices via categorical distributions. Crucially, the distribution parameter $\theta$ is now continuous and one may consider optimizing the response function
$g(\theta)=\mathbb{E}_{\iota\sim\mathrm{Ber}(\theta)}[f(\iota)]=\theta f(1)+(1-\theta)f(0)$ (7.19)
instead of $f$, having that $\min_{\theta}g(\theta)=\min_{\lambda}f(\lambda)$. Although (7.19) might look like a circular argument, the value of such reformulation is unlocked when one computes (approximate) hypergradients via appropriate continuous relaxation or discrete gradient estimators [niculae2023discrete]. These techniques may be combined with the procedures described in this chapter, making it possible to avoid the exponential cost of computing $f$ on all the configuration. This is still an active area of research: successful examples include the optimization of graph structures in transductive learning [franceschi2019learning] and topological order of nodes in DAG learning [zantedeschi2023dag], deep learning architectures [liu2018darts] and dropout coefficients [gal2017concrete].
@@ 7.7 Discussion
Implicit differentiation methods are “oblivious” to the way the approximate solution $w_{T}$ has been found. This means that they have a computational advantage in memory over the reverse-mode differentiation as they do not require storing the entire dynamics. If implemented using algorithmic differentiation, they have a run-time advantage over forward-mode differentiation since they only deal with vector quantities (the $q_{T,K}$’s). However, for this very reason, they do not directly allow computing the gradient of hyperparameters related to the optimization of the inner problem, such as the initialization mapping or the learning rate. Furthermore, they are also quite sensitive to the proprieties of the inner problem (e.g. for the fixed-point method, the dynamics must be a contraction), and therefore are applicable to a narrower class of problems. A comparative theoretical study between the three main approaches covered in this chapter is presented in [grazzi2020iteration].
Computational disadvantages of plain iterative reverse mode are related to memory requirements. Indeed, the memory complexity is $O(TW)$, with $W$ the number of weights, which can quickly become prohibitive when tuning the hyperparameters associated with modern networks. The problem has been addressed in different ways. maclaurin2015gradient suggested to completely avoid the storage of $s_{t}$, recomputing it in the backward pass by reversing the learning dynamics (7.9). Since the operation is numerically unstable, reconstructed initial weights may be very different from the original ones. As a remedy, they proposed an effective numerical strategy capable of overcoming the loss of information due to finite precision arithmetic. A general technique for reducing the memory requirements in reverse mode AD is checkpointing [Griewank08:_evaluat_deriv, Ch. 12], where only a subset of the possible checkpoints is stored, namely $w_{jC}$ for $j=1,\dots\lceil T/C\rceil$, being $C$ a given window size. In so doing, one must redo the forward pass from $jC$ to $(j+1)C-1$ to recover the missing $w_{t}$, effectively doubling the computational effort. It is easy to see that the best window size is $O(\sqrt{T})$, which allows us to complete hypergradient calculations with a memory size of $O(W\sqrt{T})$. This technique has been considered in [shaban2019truncated], together with a more radical truncation strategy similar to truncated backpropagation through time [DBLP:journals/neco/WilliamsP90], which however introduces further bias in the computation of hypergradients.
@@
In the last five chapters we covered the main classes of algorithms for hyperparamter optimization. Here, moving away from particular techniques, we discuss horizontally advanced topics in HPO such as multi-objective and constrained optimization (Section 8.1), online HPO and hyperparameter schedules (Section 8.2) and neural architecture search (Section 8.3). Next, we dedicate Section 8.4 to meta-learning, a closely related topic in the literature, discussing similarities and differences with the typical hyperparameter optimization setting. We conclude the chapter with a brief discussion on transfer of hyperparamters across model sizes, a novel topic that is relevant especially for the tuning of large foundational models (Section 8.5).
@@ 8.1 Optimizing for multiple objectives
In many real world settings there is not just one, but multiple, potentially conflicting, objectives of interest. For example, when working with hardware-constrained systems one desires predictive models that are accurate and have low resource consumption. Assume one needs to decide between two solutions $\lambda_{1},\lambda_{2}$. Each individual objective induces a distinct ordinal relationship between the candidates, but while $\lambda_{1}$ might dominate $\lambda_{2}$ in one objective, it might be worse in another. Because there is rarely a single best solution, the multi-objective hyperparameter optimization (MO) problem is to identify the Pareto front—the set of all non-dominated solutions—to help users make an informed trade-off decision. More formally, consider a function $f:\Lambda\to\mathcal{Y}$ mapping points in domain $\Lambda$ to a space $\mathcal{Y}$, which in the following we consider to be an $n$-dimensional objective space that we aim to minimize. For two points $\lambda_{1},\lambda_{2}\in\Lambda$, we say $\lambda_{1}$ weakly dominates $\lambda_{2}$ if $f(\lambda_{1})_{i}\leq f(\lambda_{2})_{i}$ for all objectives $i\in[n]$. We write $\lambda_{1}\succeq\lambda_{2}$ to denote weak domination. Furthermore, $\lambda_{1}$ is said to dominate $\lambda_{2}$ if $\lambda_{1}\succeq\lambda_{2}$ and $f(\lambda_{1})_{i}<f(\lambda_{2})_{i}$ for some objective $i$. We write $\lambda_{1}\succ\lambda_{2}$ for domination. The Pareto front $\mathcal{P}_{f}$ contains all nondominated points in $\Lambda$, formally:
$\mathcal{P}_{f}=\{\lambda\in\Lambda\,|\,\nexists\lambda^{\prime}\in\Lambda:\lambda^{\prime}\succ\lambda\}.$
Since the Pareto front is often infinite, multi-objective optimizers seek a finite approximation set $B\subset\mathcal{Y}$ of objective vectors near the Pareto front. The quality of $B$ can be measured by the dominated hypervolume indicator $\mathcal{H}$ [Zitzler1998:Multiobjective]. Given a reference point $\mathbf{r}$, $\mathcal{H}(B)$ equals the hypervolume dominated by $B$:
$\mathcal{H}(B)=\text{Vol}\left(\{\mathbf{y}\in\mathcal{Y}=\mathbb{R}^{n}\,|\,\exists\mathbf{y}^{\prime}\in B:\mathbf{y}^{\prime}\succeq\mathbf{y}\wedge\mathbf{y}\succeq\mathbf{r}\}\right).$
Computing $\mathcal{H}(B)$ involves partitioning the space into hypercubes, an operation that scales exponentially with the number of objectives $n$. This exponential scaling poses a key bottleneck for hypervolume-based multi-objective optimizers.
@@
A popular example of an MO scenario is the case of algorithmic fairness which aims to train accurate models subject to fairness criteria. Prior work in the field has proposed several fairness criteria, whose suitability depends on the application domain, and it has been shown that criteria, commonly referred to as acceptance-based and error-based, can conflict with each other [friedler2016possibility, fairness2018verma, kleinberg2018inherent]. However, obtaining solutions that satisfy various definitions to some extent and expose empirical trade-offs between various definitions can still be important from societal perspectives. Another way to tackle MO scenarios is the case of gradient-free optimization with stochastic constraints. For example, when enhancing the accuracy of a machine learning model, considerations may include constraints on model memory or prediction latency. This occurs in the practical application of deploying machine learning models on mobile devices, where memory cannot exceed a certain pre-specified threshold, or when models power devices with tight requirements on inference latency. The objective and constraint functions in such cases are real-valued and observed together. Alternatively, in tuning a deep neural network (DNN), one may aim to prevent training failures due to out-of-memory errors. Here, the constraint feedback is binary, indicating whether the constraint is violated, and the objective is not observed under constraint violation. It is common to treat constraints with equal importance to objectives, employing a surrogate model as discussed in Chapter 4. Unfeasible evaluations incur a cost, such as wasted resources or compute node failure, necessitating minimization of their occurrence. However, avoiding the unfeasible region should not compromise convergence to optimal hyperparameters. Often, optimal configurations lie at the boundary of the feasible region, presenting a challenge in balancing the costs of unfeasible evaluations and the need for exploration.